#12
Find the area of the square with sides 18 m long.
Round answers to 1 decimal place as needed and be sure to include the units.
Area =

Points possible: 2
This is attempt 2 of 3.
Score on last attempt: 0. Score in gradebook: 0
#13
Find the area of the rectangle with a length of 17 cm and a width of 16 cm.
Round answers to 1 decimal place as needed and be sure to include the units.
Area =

Points possible: 2
This is attempt 2 of 3.
Score on last attempt: 0. Score in gradebook: 0
#17
A sports ball has a diameter of 27 cm. Find the volume of the ball.
Round your answer to 2 decimal places as needed.
Volume = cm3


Points possible: 2
This is attempt 1 of 3.
Need help my friday may 02,2014 by 11:59pm thanks

Answer by rothauserc(4718)

1) A = s^2 = 18^2 = 324.0 m^2
2) A = l * w = 17 * 16 = 272.0 cm^2
3) V = (4/3)*pi*r^3 = (4/3) * pi * (13.5)^3 = 10305.99 cm^3